OpenStudy (anonymous):

can some help me plot g(X)=abs(x+2)-abs(x-3)+1

OpenStudy (mathstudent55):

Maka a table of x-y points and plot it.

OpenStudy (anonymous):

yeah,but what i am wanting is turn it into a piecewise

OpenStudy (anonymous):

theres no easy way about it, its intuitive, turning it into a peicewise function

OpenStudy (anonymous):

perhaps using a number line?

OpenStudy (anonymous):

you see in reality i have already turned it into a piecewise, expecpt i cannot get the third piece correct, so if someone could kind of explaine that.....i know i am really not making much sense lol

OpenStudy (anonymous):

technically theres nothing to do cause its not an equation.. is it = 0?

OpenStudy (anonymous):

nvm lol sry, i missted PLOT

OpenStudy (anonymous):

yes, i need plot it

OpenStudy (anonymous):

oh ok then this is prolly simplier than you realize, you just make a table and plug in x values and compute y values, then plot points

OpenStudy (anonymous):

but, can anybbody help me graph it by turing it into a piece wise adn graphing the point on the given intervals?

OpenStudy (anonymous):

or graphing the functiosn on the intervals

OpenStudy (anonymous):

what i would do, is graph and then turn it into a peicewise

OpenStudy (anonymous):

how can you do that?

OpenStudy (anonymous):

i gotta go my boss is yelling at me :/ good luck? ... y= mx +b ...

OpenStudy (anonymous):

yeah yeah, thanks tell you boss hi lol:)

OpenStudy (foolaroundmath):

If you wish to turn it into a piecewise, all you need to do is to first figure out where the "pieces" will be. Once you have done that, just evaluate the abs() function. Remember that abs(x) = x when x >= 0 and abs(x) = -x when x < 0 So for example, if you have abs(2x+3) and you want to turn it into a piecewise, first realise that the point where it "flips" is when 2x+3 = 0 i.e. x = -3/2 So, abs(2x+3) = 2x+3 when 2x+3 >= 0 or x >= -3/2 abs(2x+3) = -2x -3 when 2x+ 3 < 0 or x < -3/2

OpenStudy (mathstudent55):

I've been working on this since my first response, and this is what I came up with. The two points of interest are -2 and 3 g(x) = -(x + 2) + (x - 3) + 1 = -4, for x <= -2 g(x) = (x + 2) + (x - 3) + 1 = 2x, for -2 <= x <= 3 g(x) = (x + 2) - (x - 3) + 1 = 6, for x >= 3

OpenStudy (anonymous):

the function will be greater than zero so long as abs(x+2)>abs(x-3)+1 and less than zero while abs(x+2)<abs(x-3)+1 it just so happens that it equals zero at x=0 too

OpenStudy (anonymous):

The way to find points of interest in an absolute value function is to find when it's inner function becomes 0. That's when the sign changes and the 'absolute' part kicks in.

OpenStudy (anonymous):

So when x+2 = 0 and when x-3 = 0. You'd then solve for x.